Madison Woodworking Company bought a lathe for $75,000. It is being depreciated using MACRS with a three-year recovery period. Assuming that the lathe will lead to increased revenues of $50,000 per year, create a table showing the following items for each of years 0 through 4: before-tax cash flow; depreciation; taxable income; income tax; and after-tax cash flow. Assume a tax rate of 40%.
Answer #1

MACRS Dep rate for 3 yr property are 0.3333, 0.4445, 0.1481 & 0.0741

Tax rate = 40%

Depreciation = Purchase value * Depreciation rate

Taxable income = Net cash flow - Depreciation

Tax = Tax rate * Taxable income

ATCF = Taxable income - Tax + Depreciation

Using Excel

Year Initial cost Revenue Depreciation TI Tax ATCF
0 -75000.00 -75000.00
1 50000.00 24997.50 25002.50 10001.00 39999.00
2 50000.00 33337.50 16662.50 6665.00 43335.00
3 50000.00 11107.50 38892.50 15557.00 34443.00
4 50000.00 5557.50 44442.50 17777.00 32223.00

Showing formula in Excel

Year Initial cost Revenue Depreciation TI Tax ATCF
0 -75000 -75000
1 50000 =0.3333*75000 =C171-D171 =E171*0.4 =E171-F171+D171
2 50000 =0.4445*75000 =C172-D172 =E172*0.4 =E172-F172+D172
3 50000 =0.1481*75000 =C173-D173 =E173*0.4 =E173-F173+D173
4 50000 =0.0741*75000 =C174-D174 =E174*0.4 =E174-F174+D174